NvidiaIntely (nvidiaintely):

Find the volume of a cone with radius 30 in and slant height 50 in.

NvidiaIntely (nvidiaintely):

@Kevin

OpenStudy (welshfella):

volume of a cone = (1/3)pi r^2 h where r = radius of base and h = altitude

OpenStudy (abhinav12):

r = 30 in slant ht , l= 50in first find height h h ^2 = l^2 - r^2 h^2 = 2500 - 900 h^2 = 1600 h = 40in volume V= pi r^2 h /3 => V= pi x 900 x 40 /3 => V = 1200 pi in ^3

OpenStudy (welshfella):

You can find the altitude from the radius and the slant height using pythagogoras